OpenStudy (anonymous):

which ordered pairs from the replacement set are solutions to the equation? y=-2x+9; replacement set {(-1,11),(0,8),(1,7),(2,5)} check all that apply. a.(-1,11) b. (0,8) c.(1,7) d.(2,5)

OpenStudy (anonymous):

Basically you're setting x = -1, x = 0, x = 1, and x = 2 and seeing if you get y = 11, y= 8, y = 7 and y = 5 as the result Example: y = -2x + 9 "the y coordinate is found by taking negative two times the x coordinate and then adding nine to that." let x = -1 y = -2*(-1) + 9 ==> -2 times -1 is +2 y = 2 + 9 add 2 and 9 y = 11 So the ordered pair is (-1, 11)